Question 14 (1 point) A nice private pay nursing home in Estonia may be contracted for 1,400 Euros per month. US private pay nursing homes cost 4,500 per month. The number of Euros per dollar is 1.0 According to PPP what should a nursing home similar to the one in Estonia cost in the USA in dollars? Your Answer:

Answer Question 15 (1 point) Anderson Manufacturing has current assets of $150 million, current liabilities of $95, inventory of $90, and cash of $5. What is Anderson's Quick Ratio?
Answer #1

14). $/Euro = 1 / [Euro/$] = 1 / 1 = $1 / Euro

Cost in US $ = Nominal Cost in Euro * Exchange Rate[$/Euro]

= 1,400 euros * [$1 / Euro] = $1,400

15). Quick Ratio = [Current Assets - Inventory] / Current Liabilities

= [150 - 90] / 95 = 60 / 95 = 0.63